# Physics-Driven Local-Whole Elastic Deformation Modeling for Point Cloud Representation Learning

### v1

@@ 3.3 Physical Information Awareness
We set parameter values in $d$ based on the characteristics of the point cloud object to simulate its properties, including Lamé constants $\lambda$ and $\mu$, elastic modulus $E$, Poisson’s ratio $\nu$, and external forces $f$. These parameters together determine the deformation behavior of the object. At the same time, based on the configuration of deformation parameters $d$, we obtain the ground-truth displacement field $u$ using the finite element analysis method (FEM) [9]. Finally, based on the displacement field $\hat{u}$ predicted by the network and the ground-truth displacement field $u$ obtained through computation, we design two loss functions: the data-fidelity loss and the physics-informed loss. The data-fidelity loss guides the network to predict the displacement of each point in the deformed point cloud, explicitly modeling the relationship between local deformations and the whole geometric shape, thereby revealing the influence of local features on the whole structure. The physics-informed loss derives the external force distribution from the predicted local deformation data and compares it with the true external force distribution, ensuring that the network’s output adheres to the physical laws. Below, we provide a detailed description of these two loss functions.
@@
Data-fidelity Loss. Its objective is to guide the network in accurately predicting the displacement field $\hat{u}$ deformed tetrahedral mesh object and comparing it with the ground truth displacement field ${u}$, enabling it to precisely predict local geometric changes. In calculating elastic deformation, the finite element method is a common approach to solving elastic dynamics models in contact problems. Given $n$ node positions $x$, a finite element mass matrix $M$, and a hyper-elastic deformation energy $E(x)$, the solution to the contact problem can be obtained using the extremization method of the action:
[…]
@@
$S\left(u\right)=\int_{\Omega}{\left(\frac{1}{2}\sigma\colon\varepsilon\left(u\right)-f\cdot u\right)}d\Omega$ (4)
where ${\Omega}$ represents the volume region of the object, and $S(u)$ computes the total energy of the object’s deformation. $\frac{1}{2}\sigma\colon\varepsilon(u)$ represents the energy stored in the tetrahedral mesh due to elastic deformation. After the mesh is subjected to force, the resulting energy is calculated by the stress $\sigma$ and strain $\varepsilon$. $-f\cdot u$ represents the potential energy exerted by the external force, which is the work done by the external force under the displacement field $u$.
@@
${{\rm{{\cal L}}}_{df}}=\frac{1}{{\left|{\rm{{\cal D}}}\right|M}}\sum\limits_{i=1}^{\left|{\rm{{\cal D}}}\right|}{\sum\limits_{l=1}^{M}{{{\left\|{{u_{il}}-{{\hat{u}}_{il}}}\right\|}_{2}}}}$ (5)
@@
Physics-informed Loss. Based on the network’s prediction of the displacement of each tetrahedral mesh vertex, we derived the distribution of external forces on the tetrahedron using Hooke’s law [3] in linear elasticity analysis and the equilibrium equations [10]. First, each tetrahedron is composed of four vertices, denoted as ${{x}_{1}},{{x}_{2}},{{x}_{3}},{{x}_{4}}$ (initial coordinates) and ${{{x}^{\prime}}_{1}},{{{x}^{\prime}}_{2}},{{{x}^{\prime}}_{3}},{{{x}^{\prime}}_{4}}$ (coordinates after deformation). We construct the initial shape matrix ${X}=\left[{{x_{2}}-{x_{1}}\quad{\rm{}}{x_{3}}-{x_{1}}\quad{\rm{}}{x_{4}}-{x_{1}}}\right]$ and the shape matrix ${X^{\prime}}=\left[{{{x^{\prime}}_{2}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{3}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{4}}-{{x^{\prime}}_{1}}}\right]$ learned by the network. Then, we construct the deformation gradient $F$ based on the shape matrices $X$ and $X^{\prime}$:
@@
$\varepsilon=\frac{1}{2}\left(F+{{F}^{T}}\right)-I$ (7)
@@
$\sigma=\lambda Tr\left(\varepsilon\right)I+2\mu\varepsilon$ (8)
where $\lambda=\frac{E\nu}{\left(1+\nu\right)\left(1-2\nu\right)}$ and $\mu=\frac{E}{2\left(1+\nu\right)}$ is the Lamé constant, which is related to the elastic modulus $E$ and Poisson’s ratio $\nu$. These parameters, along with the external force $f$, determine how the object will deform.
@@
${{\rm{{\cal L}}}_{pi}}=\frac{1}{{\left|D\right|M}}\sum\limits_{i=1}^{\left|D\right|}{\sum\limits_{l=1}^{M}{{{\left\|{\nabla\cdot{\sigma_{il}}+{f_{il}}}\right\|}_{2}}}}$ (9)
where ${\nabla\cdot{\sigma_{il}}}$ represents the divergence of the stress tensor, which describes the force equilibrium within the mesh.
3.4 Overall Object
@@
4 Experiment
4.1 Pre-training
Baselines. We compare our method with the current mainstream self-supervised learning methods: cTree [37], Jigsaw3D [36], and Rotation [29] are based on pretext tasks. OcCo [44], GSPCon [13], and Tsai et al. [41] are based on the generative models. STRL [16], PoCCA [46], CLR-GAM [26], PointSmile [22], and ToThePoint [21] are based on contrast learning methods. CrossPoint [1], CrossNet [47], TCMSS [12], PointMCD [54], and Cosine Mixup [18] are based on cross-modal methods. Additionally, for clarity in the tables below, we use the abbreviations Sup, GM, PT, CL, and CM to represent supervised learning, generative models, pretext tasks, contrastive learning, and cross-modal methods, respectively.
@@ 4.1 Pre-training
Implementation Details. To facilitate a fair comparison with existing self-supervised learning methods, we follow the experimental setup proposed in [16] and use the backbone networks of PointNet [30] and DGCNN [45] as the encoder ${{\Phi}_{\theta}}$. We use Adam [17] optimizer with weight decay $1\times{{10}^{-4}}$ and initial learning rate $1\times{{10}^{-3}}$. The learning rate scheduler is cosine annealing [25], and the model is trained end-to-end across 200 epochs. After pre-training, all downstream tasks are performed on the pre-trained encoder ${{\Phi}_{\theta}}$. Our models are trained on two Tesla V100-PCI-E-32G GPUs. For the classification task, we use a pre-trained backbone model to extract point cloud features on the ModelNet40 dataset and ScanObjectNN dataset, and these features are used to train a simple linear Support Vector Machine (SVM) classifier [40]. We also evaluate the few-shot results on two classification datasets. For segmentation tasks, We evaluate part segmentation using the mean intersectionover-union (mIoU) metric, which is computed by averaging the IoU of each part in the object and then averaging the values obtained from each object category. In dataset S3DIS, we performed 6-fold cross-validation on its 6 areas to obtain the final segmentation results.
Table 1: Comparison of linear classification accuracy based on ModelNet40 dataset.
Method Category PointNet Acc DGCNN Acc
PointNet [30] Sup 89.2 --
DGCNN [45] -- 92.2
Jigsaw3D [36] PT 87.3 90.6
Rotation [29] 88.6 90.8
STRL [16] CL 88.3 90.9
ToThePoint [21] $85.6$ 89.2
CLR-GAM [26] 88.9 91.3
PoCCA [46] 89.4 91.4
PointSmile [22] 90.0 91.8
CrossPoint [1] CM 89.1 91.2
PointMCD [54] 90.0 --
TCMSS [12] 87.7 91.4
Cosine Mixup [18] 88.7 91.7
CrossNet [47] 89.5 91.5
OcCo [44] GM 88.7 89.2
GSPCon [13] 89.8 91.5
Tsai et al. [41] 90.1 92.0
Ours 90.3 92.2
4.2 Object Classification
@@
Table 2: Comparison of linear classification accuracy based on ScanObjectNN dataset.
Method Category PointNet Acc DGCNN Acc
PointNet [30] Sup 73.3 --
DGCNN [45] -- 82.8
Jigsaw3D [36] PT 55.2 59.5
STRL [16] CL 74.2 77.9
ToThePoint [21] 74.7 81.9
CLR-GAM [26] 75.7 82.1
PoCCA [46] 75.6 82.2
Pointsmile [22] 75.8 82.8
CrossPoint [1] CM 75.6 81.7
TCMSS [12] 75.1 84.0
Cosine Mixup [54] 72.5 84.7
CrossNet [47] 76.8 83.9
OcCo [44] GM 69.5 78.3
GSPCon [13] 75.8 83.9
Tsai et al. [41] 70.6 74.8
Ours 77.3 86.4
[…]
Table 3: Few-shot object classification results on ModelNet40 dataset. We report mean and standard error over 10 runs.
Method 5-way 10-way
10-shot 20-shot 10-shot 20-shot
Jigsaw3D [36] 34.3$\pm$1.3 42.2$\pm$3.5 26.0$\pm$2.4 29.9$\pm$2.6
cTree [37] 60.0$\pm$2.8 65.7$\pm$2.6 48.5$\pm$1.8 53.0$\pm$1.3
OcCo [44] 90.6$\pm$2.8 92.5$\pm$1.9 82.9$\pm$1.3 86.5$\pm$2.2
GSPCon [13] 93.6$\pm$2.9 95.3$\pm$2.5 86.1$\pm$4.7 89.9$\pm$3.7
ToThePoint [21] 92.7$\pm$4.8 95.1$\pm$3.0 87.9$\pm$4.3 91.1$\pm$3.6
Cosine Mixup [18] 92.9$\pm$3.9 95.3$\pm$2.4 87.2$\pm$2.8 90.6$\pm$3.1
CrossPoint [1] 92.5$\pm$3.0 94.9$\pm$2.1 83.6$\pm$5.3 87.9$\pm$4.2
CLR-GAM [26] 93.7$\pm$1.2 96.0$\pm$2.6 87.9$\pm$2.7 91.1$\pm$1.9
TCMSS [12] 93.1$\pm$4.2 95.4$\pm$5.3 84.5$\pm$5.1 90.5$\pm$3.5
PoCCA [46] 93.5$\pm$3.7 92.1$\pm$3.6 88.1$\pm$5.3 90.9$\pm$4.0
Ours 95.7$\pm$2.6 97.8$\pm$1.9 90.4$\pm$2.7 94.8$\pm$3.0
Table 4: Few-shot object classification results on ScanobjectNN dataset. We report mean and standard error over 10 runs.
Method 5-way 10-way
10-shot 20-shot 10-shot 20-shot
Jigsaw3D [36] 65.2$\pm$3.8 72.2$\pm$2.7 45.6$\pm$3.1 48.2$\pm$2.8
cTree [37] 68.4$\pm$3.4 71.6$\pm$2.9 42.4$\pm$2.7 43.0$\pm$3.0
OcCo [44] 72.4$\pm$1.4 77.2$\pm$1.4 57.0$\pm$1.3 61.6$\pm$1.2
GSPCon [13] 78.1$\pm$3.9 81.5$\pm$3.5 64.8$\pm$3.8 73.5$\pm$2.1
ToThePoint [21] 78.1$\pm$7.9 83.8$\pm$6.1 66.7$\pm$5.4 74.2$\pm$5.0
Cosine Mixup [18] 79.6$\pm$4.6 86.4$\pm$5.0 68.5$\pm$3.2 77.0$\pm$2.3
CrossPoint [1] 74.8$\pm$1.5 79.0$\pm$1.2 62.9$\pm$1.7 73.9$\pm$2.2
CLR-GAM [26] 80.6$\pm$1.9 86.3$\pm$2.3 67.2$\pm$1.5 76.4$\pm$2.7
TCMSS [12] 79.8$\pm$6.9 83.8$\pm$6.4 67.2$\pm$3.7 74.8$\pm$4.0
PoCCA [46] 79.9$\pm$4.7 83.5$\pm$4.2 66.0$\pm$3.2 75.1$\pm$2.7
Ours 86.4$\pm$3.7 90.3$\pm$3.8 76.1$\pm$2.1 82.5$\pm$3.3
Table 5: Performance comparison of segmentation under ShapeNetPart dataset and S3DIS dataset (DGCNN as backbone).
@@
GSPCon [13] 85.7 --
Tsai et al. [41] 85.1 --
Ours 86.0 58.9
@@
Figure 5: Self-supervised training of the DGCNN network backbone is performed on the ShapeNet dataset, and t-SNE visualization of features is tested on the ModelNet10 dataset. (a) represents the result of using only physical loss. (b) represents the result of using only the implicit field loss. (c) represents our method.
[…]
4.3 Object Segmentation
We transfer the pre-trained DGCNN backbone model to the segmentation task and conduct experimental analysis using the ShapeNetPart and S3DIS datasets, fine-tuning the model in an end-to-end manner on both datasets. We report the mean IoU (Intersection-over-Union) metric in Table 5. Specifically, we perform 6-fold cross-validation on the 6 regions of the S3DIS dataset to obtain the final segmentation results. It can be seen that our method performs well on both datasets, demonstrating that the incorporation of physical information enhances the model’s segmentation capability. Figure4 shows the segmentation results on the ShapeNetPart dataset. Our method effectively segments the fine-grained details of the parts, demonstrating superior part segmentation performance.
4.4 Ablation Study
We qualitatively analyzed the effect of adding physical information on the whole detection performance. As shown in Figure 5, t-SNE [43] draws a visualization of the inputfeatures based on the ModelNet10 dataset. It can be observed that the model guided solely by physical information (as shown in (a)) is capable of learning feature representations of objects, although with relatively coarse granularity and indistinct boundaries between different object features. Furthermore, when the physics-based model is incorporated as a constraint within the data-driven pipeline (as shown in (c)), the feature learning performance significantly improves. Compared to purely data-driven methods (as shown in (b)), this hybrid methodology achieves sharper distinctions between object features and exhibits enhanced learning capabilities for characterizing objects. For some classes (e.g., desk, sofa) have imprecise class boundaries, our method still does a better job of distinguishing them.
In addition, we conducted a qualitative analysis of the performance impact of each part on the ModelNet40 and ScanObjectNN datasets. Similar to the classification experiments, we performed pre-training for each part separately, followed by training an SVM classifier using the extracted point cloud features, as summarized in Table 6. The results demonstrate that incorporating physical information as a constraint into the model learning process enhances the detection performance, enabling the model to learn more discriminative point cloud representations.
Table 6: Classification results using linear SVM for pre-trained models (DGCNN backbone) on ModelNet40 and ScanObjectNN datasets.

### v2

@@ III-B Physical Information Awareness
Based on the configuration of deformation parameters $d$, we obtain the ground-truth displacement field $u$ using the finite element analysis method (FEM) [47]. Finally, based on the displacement field $\hat{u}$ predicted by the network and the ground-truth displacement field $u$ obtained through computation, we design two loss functions: the data-fidelity loss and the physics-informed loss. The data-fidelity loss guides the network to predict the displacement of each point in the deformed point cloud, explicitly modeling the relationship between local deformations and the whole geometric shape, thereby revealing the influence of local features on the whole structure. The physics-informed loss derives the external force distribution from the predicted local deformation data and compares it with the true external force distribution, ensuring that the network’s output adheres to the physical laws. Below, we provide a detailed description of these two loss functions.
@@
Data-fidelity Loss. Its objective is to guide the network in accurately predicting the displacement field $\hat{u}$ deformed tetrahedral mesh object and comparing it with the ground truth displacement field u{u}, enabling it to precisely predict local geometric changes. In calculating elastic deformation, the finite element method is a common approach to solving elastic dynamics models in contact problems. Given $n$ node positions $x$, a finite element mass matrix $M$, and a hyper-elastic deformation energy $E(x)$, the solution to the contact problem can be obtained using the extremization method of the action:
$S\left(x\right)=\int_{0}^{T}{\left(\frac{1}{2}{{{\dot{x}}}^{T}}M\dot{x}-E\left(x\right)+{{x}^{T}}\left({{f}_{e}}+{{f}_{d}}\right)\right)}dt,$ (3)
@@
$S\left(u\right)=\int_{\Omega}{\left(\frac{1}{2}\sigma\colon\varepsilon\left(u\right)-f\cdot u\right)}d\Omega,$ (4)
where $\Omega$} represents the volume region of the object, and $S(u)$ computes the total energy of the object’s deformation. $\frac{1}{2}\sigma\colon\varepsilon(u)$ represents the energy stored in the tetrahedral mesh due to elastic deformation. After the mesh is subjected to force, the resulting energy is calculated by the stress $\sigma$ and strain $\varepsilon$. $f\cdot u$ represents the potential energy associated with external body forces under the displacement field, which is the work done by the external force under the displacement field $u$.
@@
${{\rm{{\cal L}}}_{df}}=\frac{1}{{\left|{\rm{{\cal D}}}\right|M}}\sum\limits_{i=1}^{\left|{\rm{{\cal D}}}\right|}{\sum\limits_{l=1}^{M}{{{\left\|{{u_{il}}-{{\hat{u}}_{il}}}\right\|}_{2}}}},$ (5)
@@
Physics-informed Loss. Based on the predicted displacements of tetrahedral vertices by the network, we compute the external force residuals by deriving stress tensors from predicted displacements using Hooke’s law [49] and evaluating their divergence based on the equilibrium equation [50]. This local deformation is propagated through the entire mesh via force equilibrium constraints, enforcing consistency across connected tetrahedra and shaping the whole displacement field $\hat{u}$. First, each tetrahedron is composed of four vertices, denoted as ${{x}_{1}},{{x}_{2}},{{x}_{3}},{{x}_{4}}$ (initial coordinates) and ${{{x}^{\prime}}_{1}},{{{x}^{\prime}}_{2}},{{{x}^{\prime}}_{3}},{{{x}^{\prime}}_{4}}$ (coordinates after deformation). We construct the initial shape matrix ${X}=\left[{{x_{2}}-{x_{1}}\quad{\rm{}}{x_{3}}-{x_{1}}\quad{\rm{}}{x_{4}}-{x_{1}}}\right]$ and the shape matrix ${X^{\prime}}=\left[{{{x^{\prime}}_{2}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{3}}-{{x^{\prime}}_{1}}\quad{\rm{}}{{x^{\prime}}_{4}}-{{x^{\prime}}_{1}}}\right]$ learned by the network. Then, we construct the deformation gradient $F$ based on the shape matrices $X$ and $X^{\prime}$:
@@
$\varepsilon=\frac{1}{2}\left(F+{{F}^{T}}\right)-I,$ (7)
@@
$\sigma=\lambda Tr\left(\varepsilon\right)I+2\mu\varepsilon,$ (8)
where $\lambda=\frac{E\nu}{\left(1+\nu\right)\left(1-2\nu\right)}$ and $\mu=\frac{E}{2\left(1+\nu\right)}$ is the Lamé constant, which is related to the elastic modulus $E$ and Poisson’s ratio $\nu$. $Tr\left(\varepsilon\right)$ is the trace of the strain tensor, representing the volumetric deformation. These parameters, along with the external force $f$ , determine how the object will deform.
@@
${{\rm{{\cal L}}}_{pi}}=\frac{1}{{\left|D\right|M}}\sum\limits_{i=1}^{\left|D\right|}{\sum\limits_{l=1}^{M}{{{\left\|{\nabla\cdot{\sigma_{il}}+{f_{il}}}\right\|}_{2}}}},$ (9)
where $\nabla\cdot{\sigma_{il}}$} represents the divergence of the stress tensor, which describes the force equilibrium within the mesh. $f_{il}$ denotes the ground-truth external force at the $l$-th tetrahedron of the $i$-th sample. This loss term evaluates whether each tetrahedral element satisfies physical equilibrium, thereby serving as a physics-based constraint to guide the network toward generating reasonable predictions that comply with mechanical laws.
III-C Overall Object
@@
IV Experiment
IV-A Pre-training
Baselines. We compare our method with the current mainstream self-supervised learning methods: Jigsaw3D [31] and Rotation [36] are based on pretext tasks. CPG[51], PointMoment[35], PointSmile [17], PoCCA [33], CLR-GAM [34], ToThePoint [37], and STRL [15] are based on contrast learning methods. TCMSS [39], HyperIPC[40], MCIB[41], CrossNet [16], Cosine Mixup [42], PointMCD [43], and CrossPoint [44] are based on cross-modal methods. GDANet[52], GSPCon [11], Tsai et al. [29], and OcCo [26] are based on the generative models. Additionally, for clarity in the tables below, we use the abbreviations Sup, PT, CL, CM, and GM to represent supervised learning, pretext tasks, contrastive learning, cross-modal methods, and generative models, respectively.
@@ IV-A Pre-training
Implementation Details. To facilitate a fair comparison with existing self-supervised learning methods, we follow the experimental setup proposed in [15] and use the backbone networks of PointNet [5] and DGCNN [7] as the encoder ${\Phi}_{\theta}$}. We use Adam [58] optimizer with weight decay $1\times{{10}^{-4}}$ and initial learning rate $1\times{{10}^{-3}}$. The learning rate scheduler is cosine annealing [59], and the model is trained end-to-end across 200 epochs. After pre-training, all downstream tasks are performed on the pre-trained encoder ${\Phi}_{\theta}$}. Our models are trained on two Tesla V100-PCI-E-32G GPUs. For the classification task, we use a pre-trained backbone model to extract point cloud features on the ModelNet40 dataset and ScanObjectNN dataset, and these features are used to train a simple linear Support Vector Machine (SVM) classifier [60]. For segmentation tasks, We evaluate part segmentation using the mean intersectionover-union (mIoU) metric, which is computed by averaging the IoU of each part in the object and then averaging the values obtained from each object category. In dataset S3DIS, we performed 6-fold cross-validation on its 6 areas to obtain the final segmentation results.
TABLE I: Linear classification accuracy (%) on ModelNet40 and ScanObjectNN. PN and DG denote PointNet and DGCNN backbones respectively.
Method Category ModelNet40 ScanObjectNN
PN DG PN DG
PointNet [5] Sup 89.2 - 73.3 -
DGCNN [7] - 92.2 - 82.8
Jigsaw3D [31] PT 87.3 90.6 55.2 59.5
Rotation [36] 88.6 90.8 - -
STRL [15] CL 88.3 90.9 74.2 77.9
ToThePoint [37] 85.6 89.2 74.7 81.9
CLR-GAM [34] 88.9 91.3 75.7 82.1
PoCCA [33] 89.4 91.4 75.6 82.2
PointSmile [17] 90.0 91.8 75.8 82.8
PointMoment [35] 88.8 90.9 79.0 80.5
CrossPoint [44] CM 89.1 91.2 75.6 81.7
Cosine Mixup [42] 88.7 91.7 72.5 84.7
CrossNet [16] 89.5 91.5 76.8 83.9
HyperIPC [40] - 91.8 - 84.5
TCMSS [39] 87.7 91.4 75.1 84.0
MCIB [41] - 91.6 76.6 85.4
OcCo [26] GM 88.7 89.2 69.5 78.3
Tsai et al. [29] 90.1 92.0 70.6 74.8
GSPCon [11] 89.8 91.5 75.8 83.9
Ours 90.3 92.2 79.5 86.4
Figure 4: We use t-SNE to visualise features extracted from PointNet and DGCNN on the ModelNet10 test set. Different color represents different categories.
IV-B Object Classification
@@
TABLE II: Fine-tuning accuracy (%) of shape classification on the ModelNet40 dataset and the ScanObjectNN dataset. The self-supervised pre-trained model was used as the initial weights for the supervised learning method. PN and DG denote PointNet and DGCNN.
Method Category ModelNet40 ScanObjectNN
PN DG PN DG
PointNet [5] Sup 89.2 - 73.3 -
DGCNN [7] - 92.2 - 82.8
Jigsaw3D [31] PT 89.6 92.4 76.5 82.7
STRL [15] CL - 93.1 - -
PoCCA [33] 90.2 93.2 80.3 84.1
PointSmile [17] 90.7 93.0 - -
CPG [51] - 93.6 - -
PointMCD [43] CM 91.1 93.7 - -
CrossNet [16] - 93.4 - -
OcCo [26] GM 90.1 93.1 80.0 83.9
GDANet [52] 90.1 93.3 69.0 79.4
Ours 90.6 93.9 80.9 88.0
We evaluate the trained pre-trained weights on the ModelNet10 dataset by extracting features for each sample and performing t-SNE [61] clustering visualization, as shown in Figure 4. The left and right sides of the figure visualise the PointNet and DGCNN model features, respectively. It can be observed that our pre-trained model effectively distinguishes most samples according to their categories, resulting in clear classification boundaries.
TABLE III: Performance comparison of segmentation under ShapeNetPart dataset and S3DIS dataset (DGCNN as backbone).
@@
GSPCon [11] 85.7 -
Tsai et al. [29] 85.1 -
GSPCon [11] 85.7 -
GDANet [52] - 59.0
Ours 86.0 59.0
[…]
Figure 5: Semantic segmentation results on S3DIS (DGCNN as backbone). Different colors indicate different objects. From left to right are the top view of the ground truth, the predicted top view, the front view, and the side view.
@@
IV-C Object Segmentation
To further validate the effectiveness of the physical mechanism in point cloud representation learning, we evaluate our method on object segmentation task. This task requires the model not only to understand the global geometric structure, but also to accurately capture the relative relationships between individual parts. We transfer the pre-trained DGCNN backbone model to the segmentation task and conduct experimental analysis using the ShapeNetPart and S3DIS datasets, fine-tuning the model in an end-to-end manner on both datasets. We report the mean IoU (Intersection-over-Union) metric in Table III. Specifically, we perform 6-fold cross-validation on the 6 regions of the S3DIS dataset to obtain the final segmentation results. It can be seen that our method performs well on both datasets, demonstrating that the incorporation of physical information enhances the model’s segmentation capability. Figure 5 and Figure 6 show the segmentation visualizations on the S3DIS and the ShapeNetPart dataset dataset, respectively. It can be observed that by introducing a force-driven local-to-global modeling mechanism, our method exhibits strong generalization ability across different fine-grained point cloud segmentation tasks, demonstrating a solid understanding of local details as well as the geometric relationships and structural layouts in complex indoor scenes.
Figure 7: Visualization of elastic deformation under external forces across object categories. The first row shows the original shapes of the objects, while the second row presents the deformation results predicted by the network. The red arrows indicate the direction of the applied force, while the green dashed circles highlight regions exhibiting significant deformation.
TABLE IV: Classification results using linear SVM for pre-trained models (DGCNN backbone) on ModelNet40 and ScanObjectNN datasets.